# HUMOTO: A 4D Dataset of Mocap Human Object Interactions

### v1

@@ Abstract
We present Human Motions with Objects (HUMOTO), a high-fidelity dataset of human-object interactions for motion generation, computer vision, and robotics applications. Featuring 736 sequences (7,875 seconds at 30 fps), HUMOTO captures interactions with 63 precisely modeled objects and 72 articulated parts. Our innovations include a scene-driven LLM scripting pipeline creating complete, purposeful tasks with natural progression, and a mocap-and-camera recording setup to effectively handle occlusions. Spanning diverse activities from cooking to outdoor picnics, HUMOTO preserves both physical accuracy and logical task flow. Professional artists rigorously clean and verify each sequence, minimizing foot sliding and object penetrations. We also provide benchmarks compared to other datasets. HUMOTO’s comprehensive full-body motion and simultaneous multi-object interactions address key data-capturing challenges and provide opportunities to advance realistic human-object interaction modeling across research domains with practical applications in animation, robotics, and embodied AI systems. Project: https://jiaxin-lu.github.io/humoto/.
@@ 1 Introduction
To address this, we introduce Human Motions with Objects (HUMOTO), a new 4D animation dataset captured from real performance. HUMOTO includes 736 curated sequences totaling 7,875 seconds of motion (captured at 30 fps), featuring diverse daily activities and interactions with 63 objects comprising 72 distinct parts. Many scenes involve interactions with multiple objects, such as meal preparation with various utensils, storage organization, and room arrangement. The objects span a wide range of sizes, from small household items like utensils and tools to larger furniture pieces, all modeled based on real-world measurements. All human motions are captured with detailed body and hand movements, accompanied by text annotations.
@@ 3.1 Script Development
To address the limitations of existing datasets, where interactions often appear arbitrary or disconnected, we develop a systematic approach to create action scripts before capturing a large volume of motion data. Inspired by movie production workflows of grouping actions into scenes, we established a Scene-Driven LLM Scripting framework to automate script generation. First, we created conceptual “rooms” by logically grouping related objects from our collection. We then provided these groupings to LLMs to generate cohesive interaction sequences within contextual spaces, as illustrated in Fig. 2. This resulted in rich narratives where performers executed purposeful tasks, such as opening a drawer to retrieve an item, arranging objects on a desk, or preparing a simple meal, thereby ensuring that each motion served a clear function within a broader activity sequence. Further details of the Scene-Driven LLM Scripting process are provided in the supplemental materials (Fig. 14).
@@ 3.2 Environment and Capturing
Object tracking leveraged the dual Kinect camera setup to minimize occlusions. The FoundationPose algorithm analyzed the visual data to determine 6DoF poses for each object. To address the limitations of frame-to-frame consistency assumptions during rapid movements, we implemented a dynamic reset mechanism based on mask pixel differences, reinitializing tracking when substantial movement was detected. To further improve the tracking result, we provide object masks by employing SAM2 with strategic human annotations, ensuring tracking consistency across frames where objects might be temporarily occluded.
@@ 3.3 Data Cleaning and Annotation
Dataset Key Statistics Human Motion Object Motion Contact
Detailed Max. Objects Foot Sliding Jerk MSNR Coherence Diversity Jerk Penetration Contact State
Hand in Scene ($cm$) $\downarrow$ ($m/s^{3}$) $\downarrow$ (dB) $\rightarrow$ $\uparrow$ $\uparrow$ ($m/s^{3}$) $\downarrow$ ($cm$) $\downarrow$ Entropy $\uparrow$ Consistency $\uparrow$
BEHAVE [1] 1 4.556 4.08 5.51 0.533 0.966 10.40 0.0606 2.2915 0.0667
OMOMO [39] 1 2.130 15.10 12.37 0.619 0.978 27.40 0.0602 1.9468 0.4837
IMHD [81] 1 1.474 1.14 14.20 0.554 0.951 24.06 0.1172 2.4265 0.2411
ParaHome [35] 22 3.008 9.19 1.82 0.592 0.980 0.08 0.2167 1.0254 0.6815
HUMOTO 15 0.958 1.87 9.42 0.653 0.956 1.13 0.0068 1.4587 0.5061
Mixamo - 3.184 8.14 10.88 0.616 0.958 - - - -
Table 1: Quantitative evaluation across human-object interaction datasets. Metrics defined in Sec. A.2.1 should be interpreted holistically, as optimal values depend on specific applications. The table includes two additional statistical indicators that provide context for understanding dataset characteristics. Bold indicates best, underline indicates second-best. $\uparrow$: higher values are better, $\downarrow$: lower values are better, and $\rightarrow$: values closer to Mixamo are better.
@@ 4.1 Quantitative Evaluation
We evaluate human motion, object motion, and human-object interaction using metrics such as foot sliding, jerk, penetration, contact entropy, and state consistency. These metrics offer insights into motion quality, interaction realism, and the diversity of interaction states. Additionally, we introduce a new metric, Motion Signal-to-Noise Ratio (MSNR), to assess the quality of motion relative to noise in the dataset. MSNR evaluates motion quality using the signal-to-noise ratio (SNR) [55] of joint kinematics. Higher SNR values indicate smoother motion, though excessive smoothing may result in loss of important details. We use Mixamo, an industry-standard motion capture dataset cleaned by artists, as the baseline for human motion quality. Datasets with MSNR values closer to Mixamo’s indicate comparable motion quality. Further details and metric formulations are provided in Sec. A.2.1.
@@ 4.2.2 Perceptual Study
Participants rated randomly selected videos from HUMOTO, BEHAVE, IMHD, OMOMO, and ParaHome on a 5-point Likert scale. HUMOTO achieved the highest scores in all categories: human motion (4.79$\pm$0.49), with 82% giving maximum scores), object motion (4.88$\pm$0.36), interaction quality (4.75$\pm$0.57), and overall quality (4.78$\pm$0.43). These scores significantly outperformed all comparison datasets, with the most notable difference in interaction quality, where BEHAVE scored only 2.48$\pm$1.05 and even recent datasets like IMHD (3.94$\pm$1.04) lagged considerably.
@@
These results demonstrate the superior quality of HUMOTO in both absolute ratings and direct comparisons, particularly for interaction quality and overall performance. Details are provided in Sec. A.3.
@@ Acknowledgement
The authors would like to thank Jimei Yang for their initial discussions on script development and Nathan Carr for designing and constructing the wooden stage for motion capture. We also thank Ziwen Chen for testing the MoCap suit, and Yuan Yao, Hanwen Jiang, Sanghyun Son, Shoubin Yu for helping us installing the objects. Qixing Huang is supported by NSF Career IIS-2047677, NSF IIS-2413161, and Gifts from Adobe and Google.
@@ A.1 Statistics
Dataset # hours # subjects # objects # hands Detailed Hand body Max. Obj. in Scene setup
GRAB [58] 3.8 10 51 2 1 standing
BEHAVE [1] 4.2 8 20 - 1 portable
InterCap [28] 0.6 10 10 2 1 portable
OMOMO [39] 10.1 17 15 - 1 portable
FHPA [18] 0.9 6 26 1 1 room
HOI4D [43] 22.2 9 800 1 1 room
Chairs [31] 16.2 46 70 2 1 standing
ARCTIC [11] 1.2 10 11 2 1 standing
NeuralDome 4.6 10 23 2 1 standing
TRUMANS [33] 15 7 20 2 - (proxies) room
ParaHome [35] 8.1 38 22 2 22 room
HUMOTO 2.2 1 63 2 15 scene
Table 2: Dataset statistics (contd. from Fig. 11). We provide details on the total durations, number of subjects, objects, presence of hand and body data, maximum objects in scene, and data collection setup styles.

### v2

@@ Abstract
We present Human Motions with Objects (HUMOTO), a high-fidelity dataset of human-object interactions for motion generation, computer vision, and robotics applications. Featuring 735 sequences (7,875 seconds at 30 fps), HUMOTO captures interactions with 63 precisely modeled objects and 72 articulated parts. Our innovations include a scene-driven LLM scripting pipeline creating complete, purposeful tasks with natural progression, and a mocap-and-camera recording setup to effectively handle occlusions. Spanning diverse activities from cooking to outdoor picnics, HUMOTO preserves both physical accuracy and logical task flow. Professional artists rigorously clean and verify each sequence, minimizing foot sliding and object penetrations. We also provide benchmarks compared to other datasets. HUMOTO’s comprehensive full-body motion and simultaneous multi-object interactions address key data-capturing challenges and provide opportunities to advance realistic human-object interaction modeling across research domains with practical applications in animation, robotics, and embodied AI systems. Project Page: https://jiaxin-lu.github.io/humoto/.
@@ 1 Introduction
To address this, we introduce Human Motions with Objects (HUMOTO), a new 4D animation dataset captured from real performance. HUMOTO includes 735 curated sequences totaling 7,875 seconds of motion (captured at 30 fps), featuring diverse daily activities and interactions with 63 objects comprising 72 distinct parts. Many scenes involve interactions with multiple objects, such as meal preparation with various utensils, storage organization, and room arrangement. The objects span a wide range of sizes, from small household items like utensils and tools to larger furniture pieces, all modeled based on real-world measurements. All human motions are captured with detailed body and hand movements, accompanied by text annotations.
@@ 3.1 Script Development
To address the limitations of existing datasets, where interactions often appear arbitrary or disconnected, we develop a systematic approach to create action scripts before capturing a large volume of motion data. Inspired by movie production workflows of grouping actions into scenes, we established a Scene-Driven LLM Scripting framework to automate script generation. First, we created conceptual “rooms” by logically grouping related objects from our collection. We then provided these groupings to LLMs to generate cohesive interaction sequences within contextual spaces, as illustrated in Fig. 2. This resulted in rich narratives where performers executed purposeful tasks, such as opening a drawer to retrieve an item, arranging objects on a desk, or preparing a meal, thereby ensuring that each motion served a clear function. Further details of the Scene-Driven LLM Scripting process are provided in the supplemental materials (Fig. 14).
@@ 3.2 Environment and Capturing
Object tracking leveraged the dual Kinect camera setup to minimize occlusions. The FoundationPose [66] algorithm analyzed the visual data to determine 6DoF poses for each object. To address the limitations of frame-to-frame consistency assumptions during rapid movements, we implemented a dynamic reset mechanism based on mask pixel differences, reinitializing tracking when substantial movement was detected. To further improve the tracking result, we provide object masks by employing SAM2 with strategic human annotations, ensuring tracking consistency across frames where objects might be temporarily occluded.
@@ 3.3 Data Cleaning and Annotation
Dataset Human Motion Object Motion Contact
Foot Sliding Jerk MSNR Coherence Diversity Jerk Penetration Contact State
(cmcm) $\downarrow$ ($m/s^{3}$) $\downarrow$ (dB) $\rightarrow$ $\uparrow$ $\uparrow$ ($m/s^{3}$) $\downarrow$ (cmcm) $\downarrow$ Entropy $\uparrow$ Consistency $\uparrow$
BEHAVE [1] 4.556 4.08 5.51 0.533 0.966 10.40 0.0606 2.2915 0.0667
OMOMO [39] 2.130 15.10 12.37 0.619 0.978 27.40 0.0602 1.9468 0.4837
IMHD [82] 1.474 1.14 14.20 0.554 0.951 24.06 0.1172 2.4265 0.2411
ParaHome [35] 3.008 9.19 1.82 0.592 0.980 0.08 0.2167 1.0254 0.6815
HUMOTO 0.958 1.87 9.42 0.653 0.956 1.13 0.0068 1.4587 0.5061
Mixamo 3.184 8.14 10.88 0.616 0.958 - - - -
Table 1: Quantitative evaluation across human-object interaction datasets. Metrics defined in Appendix A.2.1 should be interpreted holistically, as optimal values depend on specific applications. The table includes two additional statistical indicators that provide context for understanding dataset characteristics. Bold indicates best, underline indicates second-best. $\uparrow$: higher values are better, $\downarrow$: lower values are better, and $\rightarrow$: values closer to Mixamo are better.
@@ 4.1 Quantitative Evaluation
We evaluate human motion, object motion, and human-object interaction using metrics such as foot sliding, jerk, penetration, contact entropy, and state consistency. These metrics offer insights into motion quality, interaction realism, and the diversity of interaction states. Additionally, we introduce a new metric, Motion Signal-to-Noise Ratio (MSNR), to assess the quality of motion relative to noise in the dataset. MSNR evaluates motion quality using the signal-to-noise ratio (SNR) [55] of joint kinematics. Higher SNR values indicate smoother motion, though excessive smoothing may result in loss of important details. We use Mixamo, an industry-standard motion capture dataset cleaned by artists, as the baseline for human motion quality. Datasets with MSNR values closer to Mixamo’s indicate comparable motion quality. Further details and metric formulations are provided in Appendix A.2.1.
@@
Dataset # hours # subj. # obj. hand body max. obj. setup
GRAB [58] 3.8 10 51 1 standing
BEHAVE [1] 4.2 8 20 1 portable
InterCap [28] 0.6 10 10 1 portable
OMOMO [39] 10.1 17 15 1 portable
FHPA [18] 0.9 6 26 1 room
HOI4D [43] 22.2 9 800 1 room
Chairs [31] 16.2 46 70 1 standing
ARCTIC [11] 1.2 10 11 1 standing
NeuralDome [77] 4.6 10 23 1 standing
TRUMANS [33] 15 7 20 - (proxies) room
ParaHome [35] 8.1 38 22 22 room
HUMOTO 2.2 1 63 15 scene
Table 2: Dataset statistics. We provide details on the total durations, number of subjects, objects, presence of hand and body data, maximum objects in scene, and data collection setup styles.
@@ 4.2.2 Perceptual Study
[…]
@@
These results demonstrate the superior quality of HUMOTO in both absolute ratings and direct comparisons, particularly for interaction quality and overall performance. Details are provided in Appendix A.3.
@@ Acknowledgement
The authors would like to thank Jimei Yang for his initial discussions on script development and Nathan Carr for designing and constructing the wooden stage for motion capture. We also thank Ziwen Chen for testing the MoCap suit, and Yuan Yao, Hanwen Jiang, Sanghyun Son, Shoubin Yu for helping us install the objects. Q. Huang acknowledges the support of NSF-2047677, NSF-2413161, NSF-2504906, NSF-2515626, and GIFTs from Adobe and Google. We would like to thank UT GenAI Center for generous support of GPU hours.